OpenStudy (anonymous):

A certain radioactive isotope placed near a Geiger counter registers 160 counts per second. 32 hours later, the counter registers 10 counts per second. What is the half-life of the isotope? A. 8 hours B. 16 hours C. 24 hours D. 32 hours E. none of the above I thought it was E but I am wrong.

OpenStudy (anonymous):

@ybarrap

OpenStudy (anonymous):

it is D. i am almost positive.

OpenStudy (shane_b):

It's going to be 8 hours. You'll have to use the half life equation to solve for t1/2: \[\Large N(t)=N_0(\frac{1}{2})^{\frac{t}{\frac{1}{2}}}\]N(t) = the quantity that still remains and has not yet decayed after a time t N0 = the initial quantity of the substance that will decay t1/2 = the half-life of the decaying quantity

OpenStudy (shane_b):

C and D don't even make sense knowing that this is an exponential decay problem :)

OpenStudy (anonymous):

oh okay, thank you so much. my mistake.

OpenStudy (shane_b):

Here's what a typical exponential decay graph would look like.